OpenStudy (anonymous):

Find the inverse function of f. f(x)= x^7

OpenStudy (anonymous):

I have no idea how to solve for this. That's all the question says.

OpenStudy (anonymous):

x^7 = x*x*x*x*x*x*x, so the inverse is the 7th root of x

OpenStudy (anonymous):

To find the inverse: Replace f(x) with y Switch x's and y's, so put x where y is and x where y is. Solve for y Replace y with f^-1(x)

OpenStudy (anonymous):

I'm probably just making it harder than it is. So the answer would be f^-1=x^-7 ?

OpenStudy (anonymous):

Nope, lets do this step by step.

OpenStudy (anonymous):

\[f(x)=x^7 \implies y=x^7\] \[x=y^7\] interchange x and y's, now solve for y. \[\large \sqrt[7]{x} = \sqrt[7]{(y^7)} \implies y = \sqrt[7]{x} \implies f^{-1}(x) = \sqrt[7]{x}\]